OpenStudy (anonymous):

Anyone like super sweet word problems? Just need to know how to set this up... A contractor has a 24pound mixture that one-fourth cement and three-fourths sand. How much of a mixture that is half cement and half sand needs to be added to produce a mixture that is one-third cement?

jimthompson5910 (jim_thompson5910):

We have 24 pounds of mix. One-fourth of that is cement. So (1/4)*24 = 6 pounds of this mix is pure cement So we have the fraction 6/24 We want to add some unknown amount of pure cement to the numerator 6 (which we'll call 'x') and add some amount of mixed cement and sand (half and half). So we're adding 2x to the denominator (note: the amounts of cement and sand we're adding is x, so we're adding a total of x+x=2x pounds of mix) Therefore, after doing this, we get 6 + x -------- 24 + 2x This is the ratio of cement to the mix. We want this ratio to be 1/3, so set it equal to 1/3 to get 6 + x 1 -------- = ----- 24 + 2x 3 So your goal is to solve that equation for x to find the amount of cement you need to add. Then double it to find the total amount of mix to add.

jimthompson5910 (jim_thompson5910):

Let me know what you get for your answer

OpenStudy (anonymous):

would the common denominator be 3(24+2x)?

jimthompson5910 (jim_thompson5910):

yes, you can also just cross multiply

jimthompson5910 (jim_thompson5910):

to get 3(6+x) = 1(24+2x)

OpenStudy (anonymous):

x=6

jimthompson5910 (jim_thompson5910):

good, so you need to add 6 pounds of cement or 2*6 = 12 pounds of total mix

jimthompson5910 (jim_thompson5910):

So your final answer is 12 pounds

OpenStudy (anonymous):

it seems im supposed to apply this to the quadratic equation somehow...

jimthompson5910 (jim_thompson5910):

what do you mean

OpenStudy (anonymous):

nevermind, this stuff is like japanese to me

jimthompson5910 (jim_thompson5910):

why do you think we need to use the quadratic equation?

jimthompson5910 (jim_thompson5910):

did it say you had to?

OpenStudy (anonymous):

no but there is a solution manual in the back of the book for all odd problems, the one we were working on is an even but the answers to the questions before and after it had 2 different answers

jimthompson5910 (jim_thompson5910):

can you give me a sample problem of where it had two solutions? I'm curious now

jimthompson5910 (jim_thompson5910):

one of the odd problems you see

OpenStudy (anonymous):

haha im so embarassed, i was looking at the wrong solutions!

jimthompson5910 (jim_thompson5910):

oh lol that's alright, you had me going for a sec though

jimthompson5910 (jim_thompson5910):

hopefully the explanation I wrote above is valid (I think it is anyway) and it makes sense to you

OpenStudy (anonymous):

thanks man, any chance you could help with one more?

jimthompson5910 (jim_thompson5910):

sure, go ahead and ask

OpenStudy (anonymous):

3x^-2+2x^-1-8=0

OpenStudy (anonymous):

\[3x ^{-2}+2x ^{-1}-8=0\]

OpenStudy (anonymous):

this looks better, haha

jimthompson5910 (jim_thompson5910):

3x^(-2)+2x^(-1)-8=0 3/(x^2)+2/(x^1)-8=0 3/(x^2)+2/(x)-8=0 Now multiply EVERY term by the LCD to clear out the fractions x^2*( 3/(x^2) )+x^2*( 2/(x) )-8x^2=0*x^2 3 + 2x - 8x^2 = 0 -8x^2 + 2x + 3 = 0 Now use the quadratic formula to solve for x: x = (-b+-sqrt(b^2-4ac))/(2a) x = (-(2)+-sqrt((2)^2-4(-8)(3)))/(2(-8)) x = (-2+-sqrt(4-(-96)))/(-16) x = (-2+-sqrt(100))/(-16) x = (-2+sqrt(100))/(-16) or x = (-2-sqrt(100))/(-16) x = (-2+10)/(-16) or x = (-2-10)/(-16) x = 8/(-16) or x = -12/(-16) x = -1/2 or x = 3/4 So the solutions are x = -1/2 or x = 3/4

OpenStudy (anonymous):

ok, that one i understand a lot better

OpenStudy (anonymous):

i appreciate it

jimthompson5910 (jim_thompson5910):

yw